Sam wishes to fertilize his lawn that is 16 yards by 70 yards. The organic fertilizer costs $5.75 per bag. Each bag covers 525 square feet. What is the cost to fertilize his yard?

Answer by rfer(16322)

convert yds to ft first then to sq ft.
(16*3)(70*3)=10080 sq ft
10089/525=19.2 bags
You can't buy a part of a bad, so you buy 20 bags
20*5.75=$115.00